OpenStudy (anonymous):

Two numbers differ by 57. There sum is 185. Find the numbers. Choices for the smaller number: 54,64,74

OpenStudy (theeric):

Hello! Do you have any ideas to start with?

OpenStudy (theeric):

Are you there, @dannyjjb ?

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

um\[2n=185 \]

OpenStudy (theeric):

It's a good try! But what you're saying there is that two times a number (n) is 185. That's not what you want, though. There are two numbers, right? So, since you don't know them, you'll need two variables. Right?

OpenStudy (anonymous):

\[s +b =185 \]\[s =b-57\]

OpenStudy (theeric):

That's right! Now, do you know of any way to solve for one variable?

OpenStudy (anonymous):

I think so

OpenStudy (theeric):

Okay! :)

OpenStudy (anonymous):

\[s+s+57= 185\]\[2s =185-57\]

OpenStudy (anonymous):

s=64

OpenStudy (anonymous):

b=128

OpenStudy (triciaal):

check your b again

OpenStudy (anonymous):

k

OpenStudy (anonymous):

b=121

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

thanks everyone

OpenStudy (theeric):

Glad you got it :) You're welcome!